OpenStudy (whatever):

a 2 digit number can be written in the form 10x+y.if the digits are reserved, the new number is 10y+x.to solve coin problems, you can write one equation for the value of the coins. a collection of dimes and quarters iss worth $25. there are 40 more domes than quarters how many of each are there?

OpenStudy (anonymous):

x: dimes y: quarters \[0.1x+0.25y=25\] and \[y+40=x\]

OpenStudy (dumbcow):

D = 40 + Q 10D + 25Q = 2500 .. same thing but you can multiply equation by 100 to eliminate decimals

OpenStudy (anonymous):

solve that and you get \[x=100, y=60\]

OpenStudy (whatever):

thank you(:

OpenStudy (whatever):

how did you get that answer? when i did it i got something else